john is 4 years older than anna and joe is 2 times older than anna their ages combined is annas age times 8
Answer by thejackal(72)

let anna's age be = x
Johns age is thus = x + 4
Joes age is thus = 2(x)
Their combined age = 8x = x + x + 4 + 2x
solve for x
4x + 4 = 8x